# SPIRIT: Short-term Prediction of solar IRradIance for zero-shot Transfer learning using Foundation Models

### v1

@@ 6.1. Zero-shot Transfer Learning
Tables 1 and 2 present the results for zero-shot transfer learning, demonstrating that our model consistently outperforms the state-of-the-art baseline across both cross-location and cross-setup scenarios in both nowcasting and forecasting tasks. When transitioning between camera setups within the same location, our model consistently shows better performance relative to the baseline. However, more notably, when moving across different locations, our model achieves up to 45% improvement. In this more complex cross-location setting, our model significantly outperforms the baseline, highlighting its superior generalizability and robustness. Furthermore, even in the traditional setup where models are trained and tested on the same location, our approach demonstrates enhanced forecasting performance, further emphasizing its effectiveness across diverse deployment conditions.
6.2. Fine-tuning with Limited Data
@@
For the analysis of fine-tuning results, we merge cross-setup and cross-location scenarios to ensure a sufficient number of data points for robust confidence interval plots, as depicted in both the nowcasting (Figure 2) and forecasting (Figure 3) tasks. Since nowcasting is a relatively simpler task, both models exhibit rapid improvement within the first week. However, the baseline model reaches performance saturation early, at approximately 45%, while our model continues to reduce its error, achieving a significant improvement, dropping below 20% within four weeks.
@@
In forecasting, SPIRIT consistently outperforms the baseline, demonstrating notably lower variance, particularly in data-limited settings (0-2 weeks of data). This underscores SPIRIT’s superior stability and reliability, with its nMAP error remaining consistently below that of the baseline. In contrast, the baseline model exhibits higher variance, indicating greater inconsistency and confusion in its performance when limited data is available. Both models experience a typical performance decline as the forecasting horizon extends from 1-hour to 4-hour forecasts, driven by the increased uncertainty over longer time horizons. Nonetheless, SPIRIT’s consistently lower variance and sustained performance highlight its robustness and its ability to adapt more effectively to challenging conditions. The transition from a zero-shot configuration to fine-tuning results in noticeable performance improvements; however, the gains diminish after approximately eight weeks of fine-tuning, suggesting that extended fine-tuning beyond this period yields only marginal additional benefits. All the results are in Appendix C.
7. Ablation Studies
7.1. Investigating Different Vision Encoders
We examine the impact of different vision models on SPIRIT’s performance, also highlighting the versatility of our system across different foundation models. We evaluate the CNN-based ResNet-152 (He et al., 2016), the vision transformer-based DINOv2 Giant (Oquab et al., 2023), and our implementation using Google ViT-Huge (Dosovitskiy, 2020). Results summarized in Table 3, demonstrate that the ViT-based models consistently outperform the ResNet-152 CNN model, which can be attributed to the superior capability of ViT architectures in capturing global image features (Jeeveswaran et al., 2022).
Table 3. We explore the impact of using different vision encoders on the overall model performance for nowcasting and forecasting, with training on TSI 2020 and testing on TSI 2021, measured by nMAP error.
Model Nowcast Forecast
@@ 7.1. Investigating Different Vision Encoders
ResNet-152 10.50 24.56 27.82 31.23 35.85
DINOv2 Giant 9.74 21.22 23.56 27.93 33.13
Google ViT-Huge 9.32 19.96 22.64 26.30 31.58
7.2. Foundation Model Size
Table 4 presents an analysis of how the size of the foundation model influences the performance of our nowcasting and forecasting architectures. Although increasing model size has traditionally been linked to performance gains, we observe that beyond a certain threshold, further scaling yields diminishing returns. This suggests that larger models do not always lead to better performance. In fact, models with 304M and 86M parameters outperform their larger counterparts with 632M parameters in forecasting and nowcasting, respectively. This aligns with recent work, which highlights that adjusting model size based on a computational budget, rather than blindly increasing model size, can lead to more efficient architectures with reduced inference costs (Alabdulmohsin et al., 2023).
Table 4. We evaluate the impact of varying size of the Google ViT vision encoder on the overall performance of the model for both nowcasting and forecasting tasks, with training on TSI 2020 and testing performed on TSI 2021.
Model Parameters Nowcast Forecast
@@ 7.2. Foundation Model Size
86M 9.14 21.92 24.07 28.73 34.50
@@
632M 9.32 19.96 22.64 26.30 31.58
8. Conclusion
This work addresses a critical challenge in solar irradiance forecasting: adapting models to new geographic locations with no prior data. By utilizing transfer learning and pre-trained models, SPIRIT generalizes well to new locations, reducing the reliance on large, location-specific datasets. As more site-specific data becomes available post-deployment, the system can be effectively fine-tuned, improving prediction accuracy and supporting better energy yield estimates and operational planning. Additionally, SPIRIT’s modular design allows for the seamless integration of any emerging vision models, ensuring that the framework remains up-to-date with the latest advancements. This scalable solution for solar irradiance forecasting can accelerate the deployment of solar farms—particularly in remote and emerging markets. SPIRIT supports the transition to renewable energy by enhancing the reliability, cost-effectiveness, and accessibility of solar energy generation.
9. Future Work and Limitations
One key limitation is that the datasets used for evaluation are all from North America, largely due to the limited availability of publicly accessible datasets from other regions. Specifically, the solar movement patterns and dynamics change in the Southern Hemisphere and need to be studied. To improve the generalizability of our system, future work will incorporate data from other continents. Additionally, while our model performs well, the use of foundation models introduces real-time inference costs and computational overheads. Future efforts will focus on reducing computational efficiency, enabling deployment on resource-constrained edge devices without sacrificing accuracy.
Appendix A Dataset Details
Table 5. A Comparative Overview of the TSI880, ASI16, and SKIPP’D Datasets: Key Attributes Including Geographical Location, Data Provided, Image Resolution, Collection Frequency, and Annual Sample Size
@@ Appendix A Dataset Details
A.1. Overview of Datasets
@@
Figure 4. Examples of sunny, partly cloudy, and overcast conditions, captured by different sky cameras, are shown from left to right, across the three datasets: TSI, ASI, and SKIPP’D, displayed from top to bottom.
A.2. Temporal Consistency in Forecasting
@@
Appendix B Clear Sky Global Horizontal Irradiance
@@
Clear Sky GHI is computed using mathematical models incorporating solar position, atmospheric transmittance, and radiative transfer principles. A common approach is the Ineichen-Perez model (Stein et al., 2012):
(5) $GHI_{\text{clear}}=I_{0}\cdot\tau\cdot\cos(\theta_{z})$
@@
B.1. Extraterrestrial Irradiance ($I_{0}$)
Extraterrestrial irradiance ($I_{0}$) is the solar irradiance just outside Earth’s atmosphere, slightly varying due to Earth’s elliptical orbit around the Sun. It is given by:
(6) $I_{0}=S_{c}\cdot\left(1+0.033\cos\left(\frac{2\pi n}{365}\right)\right)$
where $S_{c}=1367$ W/m² (solar constant), $n$ is the day of the year (1 for January 1, 365 for December 31).
B.2. Atmospheric Transmittance ($\tau$)
The atmospheric transmittance $\tau$ accounts for the attenuation of solar radiation by the atmosphere. It is often estimated using empirical models, such as the Ineichen-Perez model (Stein et al., 2012):
(7) $\tau=a\cdot e^{-b\cdot m}$
where $a,b$ are empirical coefficients dependent on location and aerosol content, $m$ is the air mass, given by (Ineichen and Perez, 2002):
(8) $m=\frac{1}{\cos(\theta_{z})+0.15(93.885-\theta_{z})^{-1.253}}$
where $\theta_{z}$ is the solar zenith angle.
Appendix C Fine-tuning Detailed Results
C.1. Nowcasting
@@
Detailed results for these experiments are presented in Tables 6-9.
Table 6. Nowcasting Performance with 1 week training
Trained on Finetuned on SPIRIT Gao et al. (Gao and Liu, 2022)
@@
Table 7. Nowcasting Performance with 2 weeks training
Trained on Finetuned on SPIRIT Gao et al. (Gao and Liu, 2022)
@@
Table 8. Nowcasting Performance with 3 weeks training
Trained on Finetuned on SPIRIT Gao et al. (Gao and Liu, 2022)
@@
Table 9. Nowcasting Performance with 4 weeks training
Trained on Finetuned on SPIRIT Gao et al. (Gao and Liu, 2022)
@@
C.2. Forecasting
Table 10. Forecasting Performance with 2 weeks of training.
Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
@@ C.2. Forecasting
Table 11. Forecasting Performance with 4 weeks of training.
Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
@@
Table 12. Forecasting Performance with 8 weeks of training.
Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
@@
Table 13. Forecasting Performance with 12 weeks of training.
Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
@@
Table 14. Forecasting Performance with 16 weeks of training.
Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
@@
The results are presented in Tables 10, 11, 12, 13, and 14. Figure 3 was constructed by systematically aggregating the results from our fine-tuning experiments, encapsulating the performance trends observed across different training durations. By leveraging visualization techniques, the figure provides a holistic representation of how the model adapts as more site-specific data becomes available. It effectively summarizes variations in performance across different random splits of training data and across different sets of source and target datasets.

### v2

@@ 5.1 Zero-shot Transfer Learning
Tables 1 and 2 present the results for zero-shot transfer learning, demonstrating that our model consistently outperforms the state-of-the-art baseline across both cross-location and cross-setup scenarios in nowcasting and forecasting tasks. In nowcasting, SPIRIT exhibits remarkable adaptability when transitioning between camera setups within the same location (TSI to ASI and vice versa) or across different geographical contexts (TSI and ASI to SKIPP’D), outperforming the baseline by up to 70%. Similarly, in forecasting tasks, our approach achieves up to 45% improvement when transferring across locations (from TSI/ASI to SKIPP’D). These substantial performance gains, particularly in challenging cross-location and cross-task settings, underscore our model’s exceptional generalizability and robustness. Furthermore, even in the traditional setup where models are trained and tested on the same dataset and setup, SPIRIT demonstrates enhanced or comparable forecasting performance, further emphasizing its effectiveness across diverse deployment conditions. Additional results for our system with different foundation model variants are provided in Appendix E, confirming that the observed improvements stem from SPIRIT’s design.
5.2 Fine-tuning with Limited Data
@@
Our experimental analysis merges cross-setup evaluations (training on ASI, finetuning and subsequently testing on TSI, and vice versa) and cross-location scenarios (shifting from ASI or TSI to SKIPP’D) to ensure a sufficient number of experimental instances for the robust confidence interval plots in Figures 2 and 3 for nowcasting and forecasting respectively. In the simpler nowcasting task, both models exhibit rapid improvement within the first week. However, the baseline reaches performance saturation around 45%, while SPIRIT continues to reduce its error, achieving a significant improvement, dropping below 20% within four weeks.
@@
In forecasting, SPIRIT consistently outperforms the baseline, demonstrating notably lower variance, particularly in data-limited settings in the first 2 weeks. This underscores SPIRIT’s superior stability and reliability, with its nMAP error remaining consistently below that of the baseline. In contrast, the baseline exhibits higher variance, indicating greater inconsistency and confusion in its performance when limited data is available. Both models experience a typical performance decline as the forecasting horizon extends from 1-hour to 4-hour forecasts, driven by the increased uncertainty over longer time horizons. Nonetheless, SPIRIT’s consistently lower variance and sustained performance highlight its robustness and its ability to adapt more effectively to challenging conditions. The transition from a zero-shot configuration to fine-tuning results in noticeable performance improvements; however, the gains diminish after approximately eight weeks of fine-tuning, suggesting that extended fine-tuning beyond this period yields only marginal additional benefits. Detailed results are in Appendix D.
6 Conclusion
This work addresses a critical challenge in solar irradiance forecasting: generalizing models to novel geographic locations without historical data. SPIRIT leverages pre-trained models and physics-informed features for future covariates, enabling robust performance in data-scarce environments while facilitating post-deployment fine-tuning as site-specific observations accumulate. SPIRIT’s modular architecture accommodates emerging vision models, ensuring its methodological relevance while its scalability accelerates the deployment of solar farms in remote and emerging markets, while also enhancing the reliability, cost-effectiveness, and accessibility of solar energy generation to facilitate the global transition toward renewable energy systems.
7 Future Work and Limitations
A key limitation is the geographical restriction of datasets to North America due to limited availability of publicly accessible data from other regions, while Southern Hemisphere solar dynamics require further investigation. Future work will expand dataset diversity to enhance generalizability. Additionally, foundation models introduce computational overhead and inference latency. Subsequent research will focus on efficiency optimization for deployment on resource-constrained edge devices.
Acknowledgements
We would like to thank the PreCog Lab at IIIT Hyderabad for their valuable guidance and discussions throughout this work. In particular, we thank Sreeram Vennam, Jain Hemang Ashok, Vaishnavi Shivkumar, Akshit Sinha, Anish Joishy and Vamshi Krishna Bonagiri for their detailed feedback and thoughtful suggestions during multiple stages of the project. We also thank Harpreet Singh, Quentin Paletta, Talha Ahmad Siddiqui, and Tejas Cavale for their valuable inputs and constructive perspectives. Finally, we thank the anonymous reviewers for their insightful feedback. The views expressed are those of the authors and do not necessarily reflect the official policies or positions of the supporting organizations.
Appendix A Ablation Studies
A.1 Impact of the Future Covariates
We investigate the contribution of the future covariates on SPIRIT’s predictive capacity, by removing them and constraining the model to rely exclusively on historical feature vectors. As demonstrated in Table 3, this modification induces performance degradation across all forecasting horizons. Error increases of 0.62, 1.01, 1.46, and 2.21 percentage points as the prediction window extends from 1-hour to 4-hour forecasts reveal a clear monotonic trend. This progressive deterioration in performance indicates that future covariates provide increasingly valuable information as the prediction horizon extends. The pronounced impact on longer-range forecasts demonstrates that future covariates become particularly crucial when predicting under heightened uncertainty, where historical patterns alone prove insufficient for accurate extrapolation. Incorporation of future covariates enables the architecture to establish more sophisticated temporal dependencies by providing complementary contextual information that augments the historical patterns encoded in past latent representations, thus enhancing the model’s ability to capture complex dynamics in the underlying time series. More details are in Section H.
Model Configuration Forecast
@@ A.1 Impact of the Future Covariates
With Future Covariates 18.96 21.77 25.46 29.89
Without Future Covariates 19.58 22.78 26.92 32.10
Table 3: Ablation study comparing the forecasting performance of SPIRIT with and without future covariates across different time horizons. We train on TSI 2020 and evaluate on TSI 2021 using nMAP error, where lower values indicate better performance.
A.2 Investigating Different Vision Encoders
We evaluate SPIRIT’s performance with diverse vision architectures, including CNN-based (ResNet, EfficientNet, ConvNextv2 (convnext)) and attention-based approaches (DINOv2, Google ViT (google_vit), Apple AIMv2 (aimv2)). Results in Table 4 show that transformer architectures yield superior performance compared to traditional convolutional approaches, corroborating findings that self-attention mechanisms better capture the global image features (vits_gt_cnns1) critical for solar forecasting. Notably, ConvNeXtv2 matches ViTs in performance, owing to its novel FCMAE framework and GRN mechanism (convnext). The architectural modularity of SPIRIT facilitates seamless integration of such emerging models, conferring substantial advantages for sustained operational deployment.
Model Params Nowcast Forecast
@@ A.2 Investigating Different Vision Encoders
AIMv2 2.72B 9.04 20.89 23.90 27.22 32.71
DINOv2 1.14B 9.74 21.22 23.56 27.93 33.13
Google 632M 9.04 18.96 21.77 25.46 29.89
ConvNeXtv2 88.7M 9.92 19.73 22.09 25.17 29.56
EfficientNet-B7 66M 13.34 25.20 27.25 29.58 34.27
ResNet-152 60.3M 10.50 24.56 27.82 31.23 35.85
Table 4: We use the largest variants of diverse CNN and transformer-based vision encoders to explore their influence on SPIRIT’s nowcasting and forecasting performance, training on TSI 2020 and testing on TSI 2021, measured by nMAP error.
A.3 Foundation Model Size
Table 5 presents an analysis of how the size of the foundation model influences the performance of our nowcasting and forecasting architectures. Although increasing model size has traditionally been linked to performance gains, we observe that beyond a threshold, further scaling yields diminishing returns or even performance degradation as with DINOv2. For transformer-based architectures, models in the moderate parameter range outperform the largest variants, while in the ResNet family, the smallest variant demonstrates superior forecasting performance, whereas the second smallest attains peak nowcasting results.
These patterns suggest that the impact of model scaling aligns more closely with parameter count than with architectural family distinctions. Specifically, for transformer-based models such as DINO and Google-ViT, performance improves consistently with increasing size up to approximately 600M parameters, beyond which gains plateau or regress. In contrast, ResNets exhibit distinct scaling dynamics that diverge from those of transformer architectures.
Model Params Nowcast Forecast
+1hr +2hr +3hr +4hr
Google ViT 86M 9.14 21.92 24.07 28.73 34.50
@@ A.3 Foundation Model Size
632M 9.04 18.96 21.77 25.46 29.89
DINOv2 22.1M 10.13 20.08 22.53 25.69 29.86
86.6M 9.56 19.52 21.95 25.00 29.57
304M 10.04 19.32 21.74 25.29 30.04
1.14B 9.74 21.22 23.56 27.93 33.13
ResNet 11.7M 10.47 22.31 24.28 27.28 32.52
25.6M 10.29 24.36 25.93 29.07 34.10
44.5M 10.97 24.99 26.21 30.62 34.86
60.3M 10.50 24.56 27.82 31.23 35.85
Table 5: Impact of scaling across Google-ViT, DINOv2, and ResNet model families on nowcasting and forecasting, with training on TSI 2020 and testing on TSI 2021, and nMAP as the error metric. Results are highlighted for the best performer in each family.
A.4 Contribution of Auxiliary and Physics Features
To quantify the individual and combined contributions of different feature modalities, we conduct a systematic ablation study examining the impact of image embeddings, physics-based features, and auxiliary spatiotemporal variables on nowcasting performance. As presented in Table 6, the results reveal several critical insights regarding feature importance and complementarity. The baseline configuration using only image embeddings achieves a test nMAP of 12.45%, establishing the fundamental predictive capability of visual information. The incorporation of either physics features (9.48% nMAP) or auxiliary features (9.55% nMAP) individually provides substantial performance improvements, with both modalities contributing nearly equivalent error reductions. Crucially, the combination of physics and auxiliary features without image embeddings yields severely degraded performance (34.23% nMAP), conclusively demonstrating that visual information serves as the primary predictive signal, while physics-based and auxiliary features provide essential complementary enhancements rather than standalone predictive power. The optimal configuration integrating all three feature types achieves the lowest error at 9.04% nMAP, validating our feature choice and the importance of our multimodal approach in capturing the complex dynamics inherent in this task.
Setup nMAP (%)
Only image embeddings 12.45
Embeddings + Physics 9.48
Embeddings + Auxiliary 9.55
Physics + Auxiliary 34.23
Embeddings + Physics + Auxiliary 9.04
Table 6: Ablation on the contribution of image embeddings, physics features, and auxiliary features for nowcasting. Results are reported as test nMAP (%, lower is better). The model is trained on TSI-2020 and tested on TSI-2021.
A.5 Variation of Regressors for Nowcasting
To establish the optimal modeling approach for nowcasting, we conducted a comprehensive evaluation of different regression models in our system while maintaining consistent input features across all models. As demonstrated in Table 7, XGBoost achieves superior performance with a test nMAP of 9.04%, substantially outperforming alternative approaches. This performance advantage can be attributed to XGBoost’s inherent strengths in handling heterogeneous tabular data, where its gradient boosting framework effectively captures complex interactions between auxiliary meteorological features, physics-based variables, and high-dimensional image embeddings. The ensemble nature of XGBoost enables it to leverage the complementary information from these diverse feature modalities more effectively than traditional neural architectures in this specific nowcasting context.
Regressor nMAP (%)
MLP 19.74
KNN Regressor 16.51
Random Forest Regressor 15.66
XGBoost 9.04
Table 7: Comparison of different regressors for nowcasting using the same input features. Results are reported as test nMAP (%, lower is better). XGBoost performs the best, indicating its suitability for this task. The models are trained on TSI 2020 and evaluated on TSI 2021.
Appendix B Dataset Details
@@
Table 8: A Comparative Overview of the TSI880, ASI16, and SKIPP’D Datasets: Key Attributes Including Geographical Location, Data Provided, Image Resolution, Collection Frequency, and Annual Sample Size
B.1 Overview of Datasets
@@
B.2 Robustness to Low-Quality Visual Inputs
The SKIPP’D dataset helps us assess the model’s robustness under challenging visual conditions. With a resolution of just 64×64 pixels, the SKIPP’D images are significantly lower in quality than those found in most practical deployments, where sky imagers typically produce higher-resolution outputs. By evaluating on SKIPP’D, we test whether our model can still perform reliably in scenarios where image quality is degraded due to hardware limitations, compression, or transmission artifacts. This setup provides a realistic stress test, ensuring that our model is not only accurate under ideal conditions, but also resilient in more constrained, real-world environments.
Figure 4: Examples of sunny, partly cloudy, and overcast conditions, captured by different sky cameras, are shown from left to right, across the three datasets: TSI, ASI, and SKIPP’D, displayed from top to bottom.
B.3 Temporal Consistency in Forecasting
@@
Appendix C Clear Sky Global Horizontal Irradiance
@@
Clear Sky GHI is computed using mathematical models incorporating solar position, atmospheric transmittance, and radiative transfer principles. A common approach is the Ineichen-Perez model :
$GHI_{\text{clear}}=I_{0}\cdot\tau\cdot\cos(\theta_{z})$ (3)
@@
Appendix D Fine-tuning Detailed Results
D.1 Nowcasting
@@
Detailed results for these experiments are presented in Tables 9, 10, 11, and 12.
Trained on Finetuned on SPIRIT wacv2022
@@
Table 9: Nowcasting performance with one week of training.
Trained on Finetuned on SPIRIT wacv2022
@@
Table 10: Nowcasting Performance with 2 weeks training
Trained on Finetuned on SPIRIT wacv2022
@@
Table 11: Nowcasting Performance with 3 weeks training
Trained on Finetuned on SPIRIT wacv2022
@@
Table 12: Nowcasting Performance with 4 weeks training
D.2 Forecasting
Interval Trained on Tested on SPIRIT wacv2022
@@ D.2 Forecasting
Table 13: Forecasting Performance with 2 weeks of training.
Interval Trained on Tested on SPIRIT wacv2022
@@
Table 14: Forecasting Performance with 4 weeks of training.
Interval Trained on Tested on SPIRIT wacv2022
@@
Table 15: Forecasting Performance with 8 weeks of training.
Interval Trained on Tested on SPIRIT wacv2022
@@
Table 16: Forecasting Performance with 12 weeks of training.
Interval Trained on Tested on SPIRIT wacv2022
@@
Table 17: Forecasting Performance with 16 weeks of training.
@@
The results are presented in Tables 13, 14, 15, 16, and 17. The figure for forecasting performance upon finetuning (Figure 3 in the main paper) was constructed by systematically aggregating the results from our fine-tuning experiments, encapsulating the performance trends observed across different training durations. By leveraging visualization techniques, the figure provides a holistic representation of how the model adapts as more site-specific data becomes available. It effectively summarizes variations in performance across different random splits of finetuning data and across different sets of source and target datasets. This covers training on TSI, and testing on ASI and SKIPP’D, as well as training on ASI, and testing on TSI and SKIPP’D. This comprehensive validation methodology substantiates SPIRIT’s efficacy across diverse deployment conditions during transfer learning.